### v1

# LIR: Efficient Degradation Removal for Lightweight Image Restoration

Abstract
Recently, there have been significant advancements in Image Restoration based on CNN and transformer. However, the inherent characteristics of the Image Restoration task are often overlooked in many works. These works often focus on the basic block design and stack numerous basic blocks to the model, leading to redundant parameters and unnecessary computations and hindering the efficiency of the image restoration. In this paper, we propose a Lightweight Image Restoration network called LIR to efficiently remove degradation (blur, rain, noise, haze, etc.). A key component in LIR is the Efficient Adaptive Attention (EAA) Block, which is mainly composed of Adaptive Filters and Attention Blocks. It is capable of adaptively sharpening contours, removing degradation, and capturing global information in various image restoration scenes in an efficient and computation-friendly manner. In addition, through a simple structural design, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks. Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art networks on most benchmarks with fewer parameters and computations. It is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic. The code is available at: https://github.com/Dongqi-Fan/LIR
Figure 1: Visualization experiment of the image restoration process. (a) The paradigm of the most modern networks that incorporate both local and global residual connections. (b) The output by Restormer still exhibits visible traces of rain. (c) The output of our LIR. Which is more plain and solves the problem of residual connections carrying degradation. (d) The visualization of the three levels of the encoder and decoder based on Restormer. The top is deblurring and bottom is deraining.
1 Introduction
Image restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, denoising, dehazing, etc. In recent years, we have witnessed the great process of deep learning methods [30, 61, 53, 35, 41, 55, 71] in IR. However, many existing works focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task. This approach often leads to parameter redundancy and unnecessary computations. Instead, our goal is an efficient and lightweight network design for various IR tasks.
To address the issue of vanishing gradient and stabilize training, as shown in Fig. 1 (a), most modern networks usually incorporate local and global residual connections. However, we think that these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output. To validate our idea, we conduct visualization experiments on the deblurring and deraining tasks using Restormer [62], a transformer model with a U-Net structure. Fig. 1 illustrates the feature maps output by the encoder and decoder at three levels, as well as the input and output images of the Restormer, on the deblurring (at the top of Fig. 1 (d)) and deraining task (at the bottom of Fig. 1 (d)). Notably, at the bottom of Fig. 1 (d), it is evident that each feature map and the final output of the network (Fig. 1 (b)) all contain degradation (rain). Although less pronounced in the deblurring task, a similar phenomenon can be observed at the top of Fig. 1 (d). To overcome this issue, we use an ingenious structural design (Fig. 2) that effectively eliminates the degradation in the local and global residual connections. (Fig. 1 (c)) shows the restored image output by our LIR, which is more plain than the Restormer (Fig. 1 (b)).
Moreover, we observe another serious issue in the visualization. In the visualization of the deblurring task (at the top of Fig. 1 (d)), the feature maps generated by encoder_level2 and decoder_level2 exhibit clear contours while the other feature maps exhibit blurred contours. This issue is even worse in the deraining task (at the bottom of Fig. 1 (d)). All feature maps not only display very blurry contours but also contain a considerable amount of noise. Especially for the feature map generated by decoder_level1, it shows barely discernible contours, only with faint traces of rain. The more important is that the high-frequency information has been lost. These observations indicate that although many modern networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and fails to address the core issues of image restoration, thus resulting in poor visual quality. Based on this insight, we propose the Adaptive Filter (Fig. 3) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
Transformer-based models have demonstrated impressive performance in various tasks, both at high-level [12, 9, 33, 49] and low-level [6, 3, 55, 52, 29]. However, there exist different opinions. [11] argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs). In addition, [31, 48, 50] share a similar viewpoint that highlights the potential of MLPs beyond transformers. On the other hand, [51] suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy. Thus, excessive computations can be avoided. Furthermore, [16, 34, 8, 73] take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer. Satisfactory results are achieved and excessive computations are avoided. Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module that aims to capture global information of the features and keeps parameters and computations friendly.
In this paper, we propose a Lightweight Image Restoration network called LIR (Fig. 2) to efficiently remove degradation and adapt to various IR tasks. The key component of LIR is the novel Efficient Adaptive Attention (EAA) Block consisting of several Adaptive Filters and Attention Blocks. Thanks to the Adaptive Filter and Attention Block, the EAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information. Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and EAA Block. This subtraction allows us to obtain the degradation, which is then expanded through the Transpose operation. After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection. Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art networks in PSNR and produces better visual results that are more in line with the human aesthetic.
Our contributions can be summarized as follows:
  • Through a simple structural design, the problem that local and global residual connections bring degradation is addressed, which is overlooked by modern networks.
  • We introduce the EAA Block, which is not only able to adaptively enhance the object contour and remove degradation but also capture global information.
  • Building upon the EAA Block, we propose the LIR, a method that efficiently removes degradation and adapts to different IR tasks, achieving comparable performance to state-of-the-art networks and better visual effects with smaller parameters and computations.
Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Efficient Adaptive Attention (EAA) Blocks. Thanks to the Adaptive Filter, Patch Attention, and Channel Attention modules, the EAA Block is endowed with the ability to extract a variety of information efficiently.
2 Related Work
2.1 Image Restoration
In recent years, IR based on deep learning methods [13, 64, 57, 27, 68, 19] have shown promising results. The transformer architectures achieve adaptive spatial aggregation at the cost of heavy computation. To address this challenge, a Kernel Basis Attention (KBA) module has been proposed by KBNet [71] in the CNN. It adaptively aggregates spatial neighborhood information in a kernel-wise manner, thereby reducing computation. ADMS [40] proposes a Filter Attribution method based on Integral Gradient (FAIG) [57] to identify the contributions of the filters to remove specific degradations. By masking filters with minimal impact on degradation removal, the network’s efficiency in removing degradations is improved. IDR [64] proposes a learnable Principal Component Analysis (PCA) and views various IR tasks as multi-task learning methods to obtain network priors and remove degradations. MPRNet [61] introduces the Multi-Stage Progressive Restoration method, which is based on the U-Net architecture and follows a multi-input multi-output style. In addition, the contrastive learning paradigm is utilized by AirNet [25] to train the network to deal with different unknown degradations. Through the prompting of the contrastive learning paradigm, the latent features in the network gradually away from the degradation space and close to the clean space. Different from the above works, our emphasis is on lightweight and efficient design while ensuring good performance.
Figure 3: The comparison between our Adaptive Filter (left) and Edge-oriented Convolution Block (right). Both are reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
2.2 Vision Transformer for Image Restoration
IPT [3] is the first work that introduces ViT [12] to low-level tasks to achieve image super-resolution, denoising, and deraining. The computational complexity of the self-attention part in ViT is $O(N^{2}D)$, where $N=H$$\times$$W$ represents the length of the patch sequence and $D$ represents the dimension of the sequence. Although vision transformers are capable of capturing long-range dependencies, the self-attention component imposes a significant computing burden. The computational complexity of the self-attention increases quadratically with the input size. To mitigate this issue, TransWeather [54] introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$. Uformer [55] and StripFormer [52] proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size. To further reduce the computational complexity, [23] estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL [29] proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention. The computational complexities are reduced to $O(NlogND)$ and $O(NMD)$ respectively, where M represents the size of feature map. However, the computation required for self-attention remains unaffordable. To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings in practical applications.
3 Method
3.1 Lightweight Image Restoration
The overall structure of LIR is illustrated in Fig. 2, which primarily consists of Efficient Adaptive Attention (EAA) Blocks. Both LIR and EAA Block adopt a down-up style to reduce computation and save parameters, which downsizing the feature map in the beginning and upsizing at the end. To preserve important low-frequency information and ensure stable training, we add two global residual connections into LIR. One connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv. As mentioned in Sec. 1, these residual connections can bring degradation. To address this, we subtract the output of the fourth EAA Block from the output of the first EAA Block to obtain the corresponding degradation. We then amplify the degradation through the Transpose operation and subtract it from the residual connection to obtain a clean connection. The same approach is used to remove degradation in the local residual connection within the EAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation. The reason we adopt this method is that the feature map of the shallow layer contains more degradation compared to a deep layer. Thus, this design allows the network to dynamically perceive this difference between shallow and deep layers during the training process and inference stage. Next, we illustrate the Adaptive Filter and Attention Block in detail.
3.2 Adaptive Filter
To enhance the object contour, remove degradation, and extract useful high-frequency information, inspired by Edge-oriented Convolution Block (ECB) [69] (at the right of Fig. 3), we develop the Adaptive Filter (AF) (at the left of Fig. 3). The AF shares similarities with the ECB, but significantly outperforms it in three ways. Firstly, the ECB introduces more convolution operations during training leading to more training costs. Instead, our AF only requires 10 training parameters resulting in smaller training costs. Secondly, the AF can be applied to various IR scenes and dynamically adjust the dependence on filters by adjusting parameters $\alpha$. However, the ECB treats each branch equally and can not adapt to various IR scenes. Lastly, the AF offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
Figure 4: The flow of the proposed Patch Attention.
Each filter in the AF has its own advantages and disadvantages. We use the learnable parameter $\alpha$ to maximize their advantages and minimize their disadvantages in different IR scenarios. Specifically, the Roberts filter is used to detect the pixel changes in the features by the differential method. It is effective in extracting prominent edges but fails to detect details. Where Roberts-x and Roberts-y represent the detection in the horizontal direction and vertical direction; The Sobel filter extracts edge information by approximating first-order differentials discretely. It has some resistance to noise and a good ability to detect details. However, the Sobel filter may cause edge thinning and false detection in scenes with a mass of noise; The Laplacian filter extracts edge information by approximating second-order differentials discretely. It can detect more detailed edges and textures than Sobel but is more sensitive to noise and may result in a bilateral effect on the edges. Where Laplacian-4 and Laplacian-8 represent 4 domains and 8 domains; The high-pass filter obtains high-frequency information by subtracting the features after Gaussian blur processing from the original features. This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the AF for stages where the features have scarce noise and a clear outline.
3.3 Attention Block
Our Attention Block consists of four components: ResBlock, Patch Attention, SimpleGate, and ResCABlock. We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module. The efficacy of the gate function has been demonstrated in prior research [31, 53, 5]. Here, we simply employ the SimpleGate module from [5] to further extract dependencies between features. Finally, we leverage the ResCABlock to extract channel-wise information.
ResBlock and ResCABlock. Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk. The only difference is that the latter has a channel attention module at the tail. The Channel Attention module [70], based on Squeeze-and-Excitation (SE) [17], has gained much popularity for its ability to extract channel-wise information. Many studies [29, 6, 5, 72] have directly utilized this module in their models. Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information.
Patch Attention. In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead. Instead, we simplify the attention map computation in our PA (Fig. 4) by using just one MLP and one element-wise multiplication. Specifically, we set C, H, and W as the channel, height, and width of the feature map, and $(P_{h},P_{w})$ as patch size. The $C_{scale}$ is used to control the dimension of the channel. We start by reducing the dimension of the input feature map from $(C,H,W)$ to $(C//C_{scale},H,W)$ using a $1\times 1$ convolution. Next, we divide the feature map into non-overlapping patches and reshape it into $(HW//P_{h}P_{w},P_{h}P_{w}C//C_{scale})$. Then, we feed this reshaped feature into the MLP to obtain the attention weights for each patch $(HW//P_{h}P_{w},1,1)$. Finally, we perform element-wise multiplication between the obtained attention weights and the input feature map to capture global information. It is important to note that our attention weights have a smaller dimension, resulting in a lower computation scale to self-attention. The $(P_{h},P_{w})$ in the LIR is set to $(8,8)$, $(4,4)$, and $(2,2)$ respectively, and the $C_{scale}$ is set to 16 and 8, according to the depth of the network.
4 Experiment
4.1 Experiment Settings
We set the number of EAA Blocks in the LIR to [3,3,3,3,4], and the width to 48. During the training, we use the $L_{1}$ loss function, AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) optimizer with an initial learning rate of $3e^{-4}$, and set weight decay as $1e^{-4}$. Random image cropping and horizontal flipping are used as data augmentation strategies. Following [62], we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising. For dehazing, we set [($128^{2}$,32), ($160^{2}$,16)] at iterations [350K, 400K]. The number of parameters and computation for each method are calculated based on the input size of (3,192,192). Specifically, parameters are counted using the PyTorch, and the computation is measured using the thop package. All experiments are conducted using the official release code for each method. For methods that do not have released code or are not based on PyTorch, we have marked them with $\triangle$ in the tables.
Method Type Rain100L [58] Params [M] GFLOPs
PSNR $\downarrow$ SSIM
UMRL [59] CNN 29.18 0.923 0.98 9.31
PReNet [44] RNN 32.44 0.950 0.17 37.27
AirNet [25] CNN 34.90 0.966 8.93 169.47
MPRNet [61] CNN 36.40 0.965 20.13 960.39
SPAIR [42] CNN 36.93 0.969 $\triangle$ $\triangle$
HINet [4] CNN 37.28 0.970 88.67 95.92
MAXIM-2S [53] MLP 38.06 0.977 22.50 80.13
SFNet [7] CNN 38.21 0.974 13.27 70.00
DRCNet [26] CNN 38.23 0.976 18.90 $\triangle$
Restormer [62] Transformer 38.99 0.978 26.13 79.31
LIR (Ours) CNN 38.96 0.983 7.31 63.96
Table 1: Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
Figure 5: The visual comparisons for LIR, FFA-Net [43], GridDehazeNet [32], and AirNet [25]. LIR produces better visual results.
Figure 6: The visual comparisons for LIR, Restormer [62], SFNet [7], and HINet [4]. LIR produces state-of-the-art visual results.
4.2 Image Deraining
The Rain14000 [15] (11,200 training images) and Rain100H [58] (1,254 training images) are used to train LIR, and both the Rain100L [58] and Rain100H [58] datasets are used for evaluation. The results are presented in Tab. 1 and Tab. 2. In the Rain100L results (Tab. 1), LIR with minimal parameters and computation outperforms the DRCNet [26] and achieves comparable performance to Restormer [62] with only a slight 0.02 dB decrease. As GFLOPs do not accurately reflect the computations of the transformer model, the actual computation of LIR is much lower compared to Restormer, as shown in Tab. 1. It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet [7], the same lightweight design method as ours.
In the Rain100H results (Tab. 2), although the number of parameters and computations are reduced by 63% and 93% respectively, the PSNR is still 0.14 dB higher than MPRNet [61]. Furthermore, LIR achieves comparable performance to HINet [4]. Fig. 6 shows visual comparisons among LIR, Restormer, SFNet, and HINet. It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
Method Type Rain100H [58] Params [M] GFLOPs
PSNR $\downarrow$ SSIM
SEMI [56] CNN 16.56 0.486 $\triangle$ $\triangle$
UMRL [60] CNN 26.01 0.832 0.98 9.31
RESCAN [28] CNN 26.36 0.786 0.15 18.07
MSPFN [21] CNN 28.66 0.860 1.65 $\triangle$
MPRNet [61] CNN 30.41 0.890 20.13 960.39
HINet [4] CNN 30.65 0.894 88.67 95.92
LIR (Ours) CNN 30.58 0.904 7.31 63.96
Table 2: Image Deraining results evaluated by PSNR and SSIM on Rain100H dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
Method Type SOTS outdoor [24] Params [M] GFLOPs
PSNR $\downarrow$ SSIM
FD-GAN [10] GAN 23.15 0.920 13.98 38.90
AirNet [25] CNN 23.18 0.900 8.93 169.47
IDR [65] CNN 25.24 0.943 15.34 $\triangle$
GridDehazeNet [32] CNN 30.86 0.982 0.96 12.05
U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$
FFA-Net [43] CNN 33.38 0.980 4.46 161.74
MAXIM-2S [53] MLP 34.19 0.985 22.5 80.13
LIR (Ours) CNN 34.85 0.985 7.31 63.96
Table 3: Image dehazing results evaluated by PSNR and SSIM on SOTS outdoor dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
Figure 7: The visual comparisons for LIR, Restormer [62], AirNet [25], and MPRNet [61]. LIR produces better visual results.
4.3 Image Dehazing
The OTS [24] (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor [24] is used for evaluation. In the SOTS outdoor results (Tab. 3), LIR significantly outperforms both the MAXIM-2S [53] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB. Visual comparisons between LIR, FFA-Net, GridDehazeNet, and AirNet are shown in the Fig. 5.
Method Type CBSD68 [37] $\downarrow$ Urban100 [18] Params [M] GFLOPs
$\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
FD-GAN [10] GAN 30.25 28.81 26.43 - - - 13.98 38.90
DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$
MPRNet [61] CNN 33.54 30.89 27.56 - - - 20.13 960.39
AirNet [25] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47
BRDNet [47] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$
IDR [65] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$
DRUNet [67] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78
Restormer [5] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31
LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
Table 4: Image denoising results evaluated by PSNR and SSIM on CBSD68 and Urban100 datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch and light blue for the method that is learned to handle various noise levels.
4.4 Image Denoising
We perform denoising experiments on synthetic benchmark datasets generated by additive white Gaussian noise. The DFWB (DIV2K [1], Flickr2K, WED [36], BSD500 [2]) datasets with 8,594 training images are used to train LIR. For evaluation, we use the CBSD68 [37] and Urban100 [18] datasets. The light blue color in the Tab. 4 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50). Other methods are trained with a fixed $\sigma$ and evaluated with the same $\sigma$ as training. As shown in the Tab. 4, LIR achieves comparable performance to Restormer [62] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting. Moreover, with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [67] on Urban100. The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Fig. 7. Although LIR is not as high as Restormer on PSNR, the visual results are far superior.
4.5 Image Deblurring
The GoPro [38] (2,103 training images) dataset is used to train LIR, and the GoPro (1,111 images) and HIDE [45] (2,025 images) datasets are used for evaluation. Tab. 5 demonstrates that our LIR significantly outperforms other methods with fewer parameters and computations. Specifically, LIR is 0.13 dB higher than SPAIR [42], and 0.34 dB significantly higher than Suin et al [46].
4.6 Ablation Studies
We conduct ablation experiments on the Rain100L dataset, using unchanged LIR as the baseline. The ablation experiments are performed on three types: Simple Residual, No Patch Attention, and No Adaptive Filter. Simple Residual refers to using vanilla global and local residual connections in the model, without removing the degradation in residual connections through structural design. The results in the Tab. 6 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR. In addition, the experiment without Patch Attention demonstrates that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
Method Type GoPro [38] HIDE [45] Params [M] GFLOPs
PSNR $\downarrow$ SSIM PSNR SSIM
IDR [65] CNN 27.87 0.846 - - 15.34 $\triangle$
DBGAN [66] GAN 31.10 0.942 28.94 0.915 11.58 189.71
MT-RNN [39] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$
DMPHN [63] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$
Zhe et al. [22] CNN 31.79 0.949 - - $\triangle$ $\triangle$
Suin et al. [46] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$
SPAIR [42] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$
LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
Table 5: Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
Type PSNR Params [M] GFLOPs
Baseline 38.96 7.31 63.96
Simple Residual 38.87 6.86 49.15
No Patch Attention 38.85 6.51 62.15
No Adaptive Filter 38.66 7.31 62.40
Table 6: Ablation stuides are performed on Rain100L dataset.
5 Conclusion
In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation. Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and a simple structural design. In addition, we propose the Adaptive Filter (AF) to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks. Our AF achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Tab. 6). To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention (PA) module by leveraging an MLP and patch embedding strategy. Extensive experiments on deraining, dehazing, denoising, and deblurring demonstrate that our LIR achieves comparable performance to state-of-the-art networks with fewer parameters and computations. Notably, our approach achieves state-of-the-art visual results in deraining and denoising tasks.

### v2

@@
-LIR: Efficient Degradation Removal for Lightweight Image Restoration
+LIR: A Lightweight Baseline for Image Restoration
 Abstract
 Recently, there have been significant advancements in Image Restoration based on CNN and transformer.
-However, the inherent characteristics of the Image Restoration task are often overlooked in many works.
-These works often focus on the basic block design and stack numerous basic blocks to the model, leading to redundant parameters and unnecessary computations and hindering the efficiency of the image restoration.
-In this paper, we propose a Lightweight Image Restoration network called LIR to efficiently remove degradation (blur, rain, noise, haze, etc.).
-A key component in LIR is the Efficient Adaptive Attention (EAA) Block, which is mainly composed of Adaptive Filters and Attention Blocks.
-It is capable of adaptively sharpening contours, removing degradation, and capturing global information in various image restoration scenes in an efficient and computation-friendly manner.
-In addition, through a simple structural design, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks.
-Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art networks on most benchmarks with fewer parameters and computations.
-It is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic.
+However, the inherent characteristics of the Image Restoration task are often overlooked.
+Many works, instead, only focus on the basic block design and stack numerous such blocks to the model, leading to parameters redundant and computations unnecessary.
+Thus, the efficiency of the image restoration is hindered.
+In this paper, we propose a Lightweight Baseline for Image Restoration called LIR to efficiently reconstruct the image and remove degradations (blur, rain, noise, haze).
+First of all, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks, through a simple structural design.
+Then, to achieve lightweight, a Lightweight Adaptive Attention (LAA) Block is introduced depending on the inherent characteristics of the Image Restoration, which is mainly composed of proposed Adaptive Filters and Attention Blocks.
+LAA is capable of adaptively sharpening contours, removing degradation, and capturing global information in various Image Restoration scenes in a computation-friendly manner.
+Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art models with fewer parameters and computations in certain tasks.
+In addition, it is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic.
 The code is available at: https://github.com/Dongqi-Fan/LIR
+Keywords:
+Image Restoration Lightweight Attention
+[…]
+[width=1]figures/introduction.eps
 Figure 1: Visualization experiment of the image restoration process.
@@
 1 Introduction
-Image restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, denoising, dehazing, etc.
+Image Restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, etc.
 In recent years, we have witnessed the great process of deep learning methods [28, 59, 52, 33, 39, 54, 67] in IR.
-However, many existing works focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task.
-This approach often leads to parameter redundancy and unnecessary computations.
+However, many existing works only focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task.
+This approach often leads to parameter redundancy and computations unnecessary.
 Instead, our goal is an efficient and lightweight network design for various IR tasks.
@@
 1 illustrates the feature maps output by the encoder and decoder at three levels, as well as the input and output images of the Restormer, on the deblurring (at the top of Fig.
-1 (d)) and deraining task (at the bottom of Fig.
-1 (d)).
+1 (d)) and deraining (at the bottom of Fig.
+1 (d)) task.
 Notably, at the bottom of Fig.
@@
 1 (d).
-To overcome this issue, we use an ingenious structural design (Fig.
+To overcome this issue, we use an ingenious structural design (transpose branchs in the LIR and LAA Fig.
 2) that effectively eliminates the degradation in the local and global residual connections.
-(Fig.
-1 (c)) shows the restored image output by our LIR, which is more plain than the Restormer (Fig.
+Fig.
+1 (c) shows the restored image output by our LIR, which is more plain than the Restormer (Fig.
 1 (b)).
@@
 1 (d)), the feature maps generated by encoder_level2 and decoder_level2 exhibit clear contours while the other feature maps exhibit blurred contours.
-This issue is even worse in the deraining task (at the bottom of Fig.
-1 (d)).
+This issue is even worse in the deraining (at the bottom of Fig.
+1 (d)) task.
 All feature maps not only display very blurry contours but also contain a considerable amount of noise.
@@
 Based on this insight, we propose the Adaptive Filter (Fig.
-3) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
+3) (an component of the LAA) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
 Transformer-based models have demonstrated impressive performance in various tasks, both at high-level [12, 9, 31, 49] and low-level [6, 3, 54, 51, 27].
@@
 Satisfactory results are achieved and excessive computations are avoided.
-Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module that aims to capture global information of the features and keeps parameters and computations friendly.
+Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module (Fig.
+4) (an component of the Attention Block) that aims to capture global information of the features and keeps parameters and computations friendly.
 In this paper, we propose a Lightweight Image Restoration network called LIR (Fig.
 2) to efficiently remove degradation and adapt to various IR tasks.
-The key component of LIR is the novel Efficient Adaptive Attention (EAA) Block consisting of several Adaptive Filters and Attention Blocks.
-Thanks to the Adaptive Filter and Attention Block, the EAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information.
-Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and EAA Block.
+The key component of LIR is the novel Lightweight Adaptive Attention Block (LAA) consisting of several proposed Adaptive Filters and Attention Blocks.
+Futhermore, the Attention Block is formed by proposed Patch Attention module (Fig.
+4), ResCABlock, and ResBlock (Details can be found in Fig.
+2).
+Thanks to the Adaptive Filter and Attention Block, the LAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information.
+Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and LAA Block.
 This subtraction allows us to obtain the degradation, which is then expanded through the Transpose operation.
 After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection.
-Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art networks in PSNR and produces better visual results that are more in line with the human aesthetic.
+Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art models in PSNR and produces better visual results that are more in line with the human aesthetic.
 Our contributions can be summarized as follows:
   • Through a simple structural design, the problem that local and global residual connections bring degradation is addressed, which is overlooked by modern networks.
-  • We introduce the EAA Block, which is not only able to adaptively enhance the object contour and remove degradation but also capture global information.
-  • Building upon the EAA Block, we propose the LIR, a method that efficiently removes degradation and adapts to different IR tasks, achieving comparable performance to state-of-the-art networks and better visual effects with smaller parameters and computations.
-Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Efficient Adaptive Attention (EAA) Blocks.
-Thanks to the Adaptive Filter, Patch Attention, and Channel Attention modules, the EAA Block is endowed with the ability to extract a variety of information efficiently.
+  • Given the problem of the loss of high-frequency information observed in the visualization experiments, we propose the Adaptive Filter (Fig.
+3) as a solution to adaptively extract high-frequency information and enhance object contours.
+  • To give LIR the ability to extract global information, we propose the Patch Attention module (Fig. 4), which is more computation-friendly than the self-attention of the transformer.
+  • Based on Adaptive Filter, Patch Attention, ResCABlock, and ResBlock, we designed a lightweight and powerful block: LAA.
+It can make the LIR achieve comparable performance to state-of-the-art networks and better visual effects with smaller parameters and computations.
 2 Related Work
@@
 Different from the above works, our emphasis is on lightweight and efficient design while ensuring good performance.
-Figure 3: The comparison between our Adaptive Filter (left) and Edge-oriented Convolution Block (right).
-Both are reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
 2.2 Vision Transformer for Image Restoration
@@
 However, the computation required for self-attention remains unaffordable.
-To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings in practical applications.
+To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings.
+[…]
+[width=]figures/model_2.eps
+Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Lightweight Adaptive Attention (LAA) Blocks.
+Thanks to the Adaptive Filter, Patch Attention, ResCABlock, and the ResBlock, the LAA Block is endowed with the ability to extract a variety of information efficiently.
 3 Method
@@
 The overall structure of LIR is illustrated in Fig.
-2, which primarily consists of Efficient Adaptive Attention (EAA) Blocks.
-Both LIR and EAA Block adopt a down-up style to reduce computation and save parameters, which downsizing the feature map in the beginning and upsizing at the end.
-To preserve important low-frequency information and ensure stable training, we add two global residual connections into LIR.
+2, which primarily consists of Lightweight Adaptive Attention (LAA) Blocks.
+Both LIR and LAA Block adopt a down-up style to reduce computation and save parameters, that is downsizing the feature map in the beginning and upsizing at the end.
+To preserve low-frequency information and ensure training is stable, we add two global residual connections into LIR.
 One connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv.
@@
 1, these residual connections can bring degradation.
-To address this, we subtract the output of the fourth EAA Block from the output of the first EAA Block to obtain the corresponding degradation.
+To address this, we subtract the output of the fourth LAA Block from the output of the first LAA Block to obtain the corresponding degradation.
 We then amplify the degradation through the Transpose operation and subtract it from the residual connection to obtain a clean connection.
-The same approach is used to remove degradation in the local residual connection within the EAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation.
+The same approach is used to remove degradation in the local residual connection within the LAA Block, where we subtract the output of the fourth Attention Block from the output of the first Attention Block to obtain the corresponding degradation.
 The reason we adopt this method is that the feature map of the shallow layer contains more degradation compared to a deep layer.
 Thus, this design allows the network to dynamically perceive this difference between shallow and deep layers during the training process and inference stage.
-Next, we illustrate the Adaptive Filter and Attention Block in detail.
+Now, the problem that local and global residual connections bring degradation is addressed.
+Next, we illustrate how we can achieve lightweight and efficiency through the Adaptive Filter and Attention Block in detail.
+[…]
+[width=]figures/filters_2.eps
+Figure 3: The comparison between our Adaptive Filter (left) and Edge-oriented Convolution Block (right).
+Both are reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
 3.2 Adaptive Filter
-To enhance the object contour, remove degradation, and extract useful high-frequency information, inspired by Edge-oriented Convolution Block (ECB) [69] (at the right of Fig.
-3), we develop the Adaptive Filter (AF) (at the left of Fig.
+To achieve lightweight and efficiency, we first introduce the Adaptive Filter (at the left of Fig.
+3), inspired by Edge-oriented Convolution Block (ECB) [66] (at the right of Fig.
 3).
-The AF shares similarities with the ECB, but significantly outperforms it in three ways.
-Firstly, the ECB introduces more convolution operations during training leading to more training costs.
-Instead, our AF only requires 10 training parameters resulting in smaller training costs.
-Secondly, the AF can be applied to various IR scenes and dynamically adjust the dependence on filters by adjusting parameters $\alpha$.
+The rich filters of different kinds allow the AF to enhance the object contour, remove degradation, and extract useful high-frequency information.
+Furthermore, the LAA block can be reparameterized as one $3\times 3$ and one high-pass filter after training to achieve lightweight.
+Although Adaptive Filter shares similarities with the ECB, it significantly outperforms ECB in three ways.
+Firstly, the ECB introduces more convolution operations during training leading to more training costs (i.e.
+parameters and computations).
+Instead, our Adaptive Filter only requires 10 training parameters resulting in smaller training costs.
+Secondly, the Adaptive Filter can be applied to various IR scenes and dynamically adjust the dependence on different filters by adjusting parameters $\alpha$.
 However, the ECB treats each branch equally and can not adapt to various IR scenes.
-Lastly, the AF offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
-Figure 4: The flow of the proposed Patch Attention.
-Each filter in the AF has its own advantages and disadvantages.
-We use the learnable parameter $\alpha$ to maximize their advantages and minimize their disadvantages in different IR scenarios.
-Specifically, the Roberts filter is used to detect the pixel changes in the features by the differential method.
+Lastly, the Adaptive Filter offers a wider range of filters that can extract more types of useful information and is more robust to noisy scenes, while the ECB is brittle in rich degradation scenes.
+Each filter in the Adaptive Filter has its own advantages and disadvantages.
+Thus, we use the learnable parameters $\alpha$ to adaptively maximize their advantages and minimize their disadvantages in different LAA blocks of the LIR.
+Specifically, in these filters, the Roberts filter is used to detect the pixel changes in the features by the differential method.
 It is effective in extracting prominent edges but fails to detect details.
@@
 3.3 Attention Block
-Our Attention Block consists of four components: ResBlock, Patch Attention, SimpleGate, and ResCABlock.
+While different types of filters in Adaptive Filter (AF) enable LAA blocks to extract rich information, AF alone is not enough, as we need to take full advantage of the rich information extracted by AF.
+Thus, a strong and efficient Attention Block is introduced, which consists of four components: ResBlock, Patch Attention (PA), SimpleGate, and ResCABlock.
 We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module.
@@
 Finally, we leverage the ResCABlock to extract channel-wise information.
-ResBlock and ResCABlock.
+3.3.1 ResBlock and ResCABlock
 Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk.
-The only difference is that the latter has a channel attention module at the tail.
+The only difference between ResBlock and ResCABlock is that the latter has a channel attention module at the tail.
 The Channel Attention module [68], based on Squeeze-and-Excitation (SE) [17], has gained much popularity for its ability to extract channel-wise information.
@@
 Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information.
-Patch Attention.
+In addition, the structure of ResBlock and ResCABlock are designed to be simple to meet lightweight and efficient requirements.
+[…]
+[width=]figures/patch_attention_2.eps
+Figure 4: The flow of the proposed Patch Attention.
+3.3.2 Patch Attention
+To give the Attention Block the ability to capture global information, we introduce the Patch Attention (PA) (Fig.
+4).
 In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead.
-Instead, we simplify the attention map computation in our PA (Fig.
-4) by using just one MLP and one element-wise multiplication.
+Instead, we simplify the attention map computation in our PA by using just one MLP and one element-wise multiplication.
 Specifically, we set C, H, and W as the channel, height, and width of the feature map, and $(P_{h},P_{w})$ as patch size.
@@
 The $(P_{h},P_{w})$ in the LIR is set to $(8,8)$, $(4,4)$, and $(2,2)$ respectively, and the $C_{scale}$ is set to 16 and 8, according to the depth of the network.
+Table 1: Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
+[…]
+0.9 Method Type Rain100L [56] Params [M] GFLOPs PSNR $\downarrow$ SSIM UMRL [57] CNN 29.18 0.923 0.98 9.31 PReNet [42] RNN 32.44 0.950 0.17 37.27 AirNet [24] CNN 34.90 0.966 8.93 169.47 MPRNet [59] CNN 36.40 0.965 20.13 960.39 SPAIR [40] CNN 36.93 0.969 $\triangle$ $\triangle$ HINet [5] CNN 37.28 0.970 88.67 95.92 MAXIM-2S [52] MLP 38.06 0.977 22.50 80.13 SFNet [7] CNN 38.21 0.974 13.27 70.00 DRCNet [25] CNN 38.23 0.976 18.90 $\triangle$ Restormer [58] Transformer 38.99 0.978 26.13 79.31 LIR (Ours) CNN 38.96 0.983 7.31 63.96
+[…]
+[width=]figures/derain100l.eps
+Figure 5: The visual comparisons for LIR, Restormer [58], SFNet [7], and HINet [5]. LIR produces state-of-the-art visual results.
 4 Experiment
 4.1 Experiment Settings
-We set the number of EAA Blocks in the LIR to [3,3,3,3,4], and the width to 48.
+We set the number of LAA Blocks in the LIR to [3,3,3,3,4], and the width to 48.
 During the training, we use the $L_{1}$ loss function, AdamW ($\beta_{1}$=0.9, $\beta_{2}$=0.999) optimizer with an initial learning rate of $3e^{-4}$, and set weight decay as $1e^{-4}$.
@@
 For methods that do not have released code or are not based on PyTorch, we have marked them with $\triangle$ in the tables.
-Method Type Rain100L [58] Params [M] GFLOPs
-PSNR $\downarrow$ SSIM
-UMRL [59] CNN 29.18 0.923 0.98 9.31
-PReNet [44] RNN 32.44 0.950 0.17 37.27
-AirNet [25] CNN 34.90 0.966 8.93 169.47
-MPRNet [61] CNN 36.40 0.965 20.13 960.39
-SPAIR [42] CNN 36.93 0.969 $\triangle$ $\triangle$
-HINet [4] CNN 37.28 0.970 88.67 95.92
-MAXIM-2S [53] MLP 38.06 0.977 22.50 80.13
-SFNet [7] CNN 38.21 0.974 13.27 70.00
-DRCNet [26] CNN 38.23 0.976 18.90 $\triangle$
-Restormer [62] Transformer 38.99 0.978 26.13 79.31
-LIR (Ours) CNN 38.96 0.983 7.31 63.96
-Table 1: Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
+[…]
+[width=0.6]figures/dehaze.eps
 Figure 6: The visual comparisons for LIR, FFA-Net [41], GridDehazeNet [30], and AirNet [24]. LIR produces better visual results.
-Figure 6: The visual comparisons for LIR, Restormer [62], SFNet [7], and HINet [4]. LIR produces state-of-the-art visual results.
 4.2 Image Deraining
-The Rain14000 [15] (11,200 training images) and Rain100H [58] (1,254 training images) are used to train LIR, and both the Rain100L [58] and Rain100H [58] datasets are used for evaluation.
+The Rain14000 [15] (11,200 training images) and Rain100H [56] (1,254 training images) are used to train LIR, and the Rain100L [56] datasets is used for evaluation.
 The results are presented in Tab.
-1 and Tab.
-2.
+1.
 In the Rain100L results (Tab.
-1), LIR with minimal parameters and computation outperforms the DRCNet [26] and achieves comparable performance to Restormer [62] with only a slight 0.02 dB decrease.
+1), LIR with minimal parameters and computation outperforms the DRCNet [25] and achieves comparable performance to Restormer [58] with only a slight 0.02 dB decrease on PSNR.
+In addition, our LIR achieves state-of-the-art on SSIM, which significantly outperforms Restormer.
 As GFLOPs do not accurately reflect the computations of the transformer model, the actual computation of LIR is much lower compared to Restormer, as shown in Tab.
@@
 It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet [7], the same lightweight design method as ours.
-In the Rain100H results (Tab.
-2), although the number of parameters and computations are reduced by 63% and 93% respectively, the PSNR is still 0.14 dB higher than MPRNet [61].
-Furthermore, LIR achieves comparable performance to HINet [4].
 Fig.
@@
 It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
-Method Type Rain100H [58] Params [M] GFLOPs
-PSNR $\downarrow$ SSIM
-SEMI [56] CNN 16.56 0.486 $\triangle$ $\triangle$
-UMRL [60] CNN 26.01 0.832 0.98 9.31
-RESCAN [28] CNN 26.36 0.786 0.15 18.07
-MSPFN [21] CNN 28.66 0.860 1.65 $\triangle$
-MPRNet [61] CNN 30.41 0.890 20.13 960.39
-HINet [4] CNN 30.65 0.894 88.67 95.92
-LIR (Ours) CNN 30.58 0.904 7.31 63.96
-Table 2: Image Deraining results evaluated by PSNR and SSIM on Rain100H dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
-Method Type SOTS outdoor [24] Params [M] GFLOPs
-PSNR $\downarrow$ SSIM
-FD-GAN [10] GAN 23.15 0.920 13.98 38.90
-AirNet [25] CNN 23.18 0.900 8.93 169.47
-IDR [65] CNN 25.24 0.943 15.34 $\triangle$
-GridDehazeNet [32] CNN 30.86 0.982 0.96 12.05
-U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$
-FFA-Net [43] CNN 33.38 0.980 4.46 161.74
-MAXIM-2S [53] MLP 34.19 0.985 22.5 80.13
-LIR (Ours) CNN 34.85 0.985 7.31 63.96
 Table 2: Image dehazing results evaluated by PSNR and SSIM on SOTS outdoor dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
-Figure 7: The visual comparisons for LIR, Restormer [62], AirNet [25], and MPRNet [61]. LIR produces better visual results.
+[…]
+0.9 Method Type SOTS outdoor [23] Params [M] GFLOPs PSNR $\downarrow$ SSIM FD-GAN [11] GAN 23.15 0.920 13.98 38.90 AirNet [24] CNN 23.18 0.900 8.93 169.47 IDR [62] CNN 25.24 0.943 15.34 $\triangle$ GridDehazeNet [30] CNN 30.86 0.982 0.96 12.05 U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$ FFA-Net [41] CNN 33.38 0.980 4.46 161.74 MAXIM-2S [52] MLP 34.19 0.985 22.5 80.13 DehazeFormer [44] Transformer 34.95 0.984 2.52 13.37 LIR (Ours) CNN 34.85 0.985 7.31 63.96
+[…]
+[width=]figures/denoise.eps
+Figure 7: The visual comparisons for LIR, Restormer [58], AirNet [24], and MPRNet [59]. LIR produces state-of-the-art visual results.
 4.3 Image Dehazing
@@
 In the SOTS outdoor results (Tab.
-3), LIR significantly outperforms both the MAXIM-2S [53] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB.
-Visual comparisons between LIR, FFA-Net, GridDehazeNet, and AirNet are shown in the Fig.
-5.
-Method Type CBSD68 [37] $\downarrow$ Urban100 [18] Params [M] GFLOPs
-$\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
-FD-GAN [10] GAN 30.25 28.81 26.43 - - - 13.98 38.90
-DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$
-MPRNet [61] CNN 33.54 30.89 27.56 - - - 20.13 960.39
-AirNet [25] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47
-BRDNet [47] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$
-IDR [65] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$
-DRUNet [67] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78
-Restormer [5] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31
-LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
+2), LIR significantly outperforms both the MAXIM-2S [52] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB.
+In addition, LIR achieves comparable performance on PSNR and better performance on SSIM to DehazeFormer.
+It is worth noting that GFLOPs and parameters do not accurately reflect the complexity of the transformer model, the actual complexity of LIR is much lower compared to DehazeFormer.
+Visual comparisons are shown in the Fig.
+6.
 Table 3: Image denoising results evaluated by PSNR and SSIM on CBSD68 and Urban100 datasets.
 Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch and light blue for the method that is learned to handle various noise levels.
+[…]
+0.9 Method Type CBSD68 [35] $\downarrow$ Urban100 [18] Params [M] GFLOPs $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ FD-GAN [11] GAN 30.25 28.81 26.43 - - - 13.98 38.90 DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$ MPRNet [59] CNN 33.54 30.89 27.56 - - - 20.13 960.39 AirNet [24] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47 BRDNet [46] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$ IDR [62] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$ DRUNet [64] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78 Restormer [4] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31 LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
 4.4 Image Denoising
@@
 As shown in the Tab.
-4, LIR achieves comparable performance to Restormer [62] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting.
-Moreover, with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [67] on Urban100.
+3, LIR achieves comparable performance to Restormer [58] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting.
+Moreover, it is worth noting that with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [64] on Urban100, while the DRUNet is trained under fixed $\sigma$ setting and our LIR is under random $\sigma$ setting.
 The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Fig.
 7.
-Although LIR is not as high as Restormer on PSNR, the visual results are far superior.
+Although LIR is not as high as Restormer on PSNR, the state-of-the-art visual result is achieved.
+Table 4: Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
+[…]
+0.9 Method Type GoPro [36] HIDE [43] Params [M] GFLOPs PSNR $\downarrow$ SSIM PSNR $\downarrow$ SSIM IDR [62] CNN 27.87 0.846 - - 15.34 $\triangle$ DBGAN [65] GAN 31.10 0.942 28.94 0.915 11.58 189.71 MT-RNN [37] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$ DMPHN [60] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$ Zhe et al.
+[21] CNN 31.79 0.949 - - $\triangle$ $\triangle$ Suin et al.
+[45] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$ SPAIR [40] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$ LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
 4.5 Image Deblurring
@@
 Simple Residual refers to using vanilla global and local residual connections in the model, without removing the degradation in residual connections through structural design.
+In addition, No Adaptive Filter refers to using a vanilla convolution to replace the Adaptive Filter for fair comparison.
+Therefore, the number of parameters is unchanged.
 The results in the Tab.
@@
 In addition, the experiment without Patch Attention demonstrates that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
-Method Type GoPro [38] HIDE [45] Params [M] GFLOPs
-PSNR $\downarrow$ SSIM PSNR SSIM
-IDR [65] CNN 27.87 0.846 - - 15.34 $\triangle$
-DBGAN [66] GAN 31.10 0.942 28.94 0.915 11.58 189.71
-MT-RNN [39] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$
-DMPHN [63] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$
-Zhe et al. [22] CNN 31.79 0.949 - - $\triangle$ $\triangle$
-Suin et al. [46] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$
-SPAIR [42] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$
-LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
-Table 5: Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
-Type PSNR Params [M] GFLOPs
-Baseline 38.96 7.31 63.96
-Simple Residual 38.87 6.86 49.15
-No Patch Attention 38.85 6.51 62.15
-No Adaptive Filter 38.66 7.31 62.40
 Table 5: Ablation stuides are performed on Rain100L dataset.
-5 Conclusion
+Metric Baseline Simple Residual Patch Attention Adaptive Filter
+PSNR 38.96 38.87 38.85 38.66
+Params [M] 7.31 6.86 6.51 7.31
+GFLOPs 63.96 49.15 62.15 62.40
+5 Conclusion and Limitation
+5.1 Conclusion
 In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation.
@@
 Our AF achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Tab.
-6).
+5).
 To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention (PA) module by leveraging an MLP and patch embedding strategy.
-Extensive experiments on deraining, dehazing, denoising, and deblurring demonstrate that our LIR achieves comparable performance to state-of-the-art networks with fewer parameters and computations.
-Notably, our approach achieves state-of-the-art visual results in deraining and denoising tasks.
+5.2 Limitation
+For the deraining task, LIR achieves comparable performance with the state-of-the-art method on PSNR and the best performance on SSIM with smaller parameter numbers and computations; For denoising and dehazing tasks, LIR also achieves comparable performance to the state-of-art method.
+In addition, LIR achieves state-of-the-art visual results on deraining and denoising tasks.
+However, as Tab.
+4 shows, LIR is less prominent on deblurring tasks than deraining and denoising.
+We believe that this is because, in the deblurring task, object outlines tend to be stretched out very long.
+While our Adaptive Filter is very effective when dealing with static object contours, it has limitations when dealing with dynamic contours caused by fast-moving objects.

### v3

@@
 LIR: A Lightweight Baseline for Image Restoration
-Abstract
+Abstract.
 Recently, there have been significant advancements in Image Restoration based on CNN and transformer.
-However, the inherent characteristics of the Image Restoration task are often overlooked.
-Many works, instead, only focus on the basic block design and stack numerous such blocks to the model, leading to parameters redundant and computations unnecessary.
+However, the inherent characteristics of the Image Restoration task are often overlooked in many works.
+They, instead, tend to focus on the basic block design and stack numerous such blocks to the model, leading to parameters redundant and computations unnecessary.
 Thus, the efficiency of the image restoration is hindered.
-In this paper, we propose a Lightweight Baseline for Image Restoration called LIR to efficiently reconstruct the image and remove degradations (blur, rain, noise, haze).
-First of all, LIR addresses the degradations existing in the local and global residual connections that are ignored by modern networks, through a simple structural design.
-Then, to achieve lightweight, a Lightweight Adaptive Attention (LAA) Block is introduced depending on the inherent characteristics of the Image Restoration, which is mainly composed of proposed Adaptive Filters and Attention Blocks.
-LAA is capable of adaptively sharpening contours, removing degradation, and capturing global information in various Image Restoration scenes in a computation-friendly manner.
-Extensive experiments demonstrate that our LIR achieves comparable performance to state-of-the-art models with fewer parameters and computations in certain tasks.
-In addition, it is worth noting that our LIR produces better visual results than state-of-the-art networks that are more in line with the human aesthetic.
+In this paper, we propose a Lightweight Baseline network for Image Restoration called LIR to efficiently restore the image and remove degradations.
+First of all, through an ingenious structural design, LIR removes the degradations existing in the local and global residual connections that are ignored by modern networks.
+Then, a Lightweight Adaptive Attention (LAA) Block is introduced which is mainly composed of proposed Adaptive Filters and Attention Blocks.
+The proposed Adaptive Filter is used to adaptively extract high-frequency information and enhance object contours in various IR tasks, and Attention Block involves a novel Patch Attention module to approximate the self-attention part of the transformer.
+On the deraining task, our LIR achieves the state-of-the-art Structure Similarity Index Measure (SSIM) and comparable performance to state-of-the-art models on Peak Signal-to-Noise Ratio (PSNR).
+For denoising, dehazing, and deblurring tasks, LIR also achieves a comparable performance to state-of-the-art models with a parameter size of about 30%.
+In addition, it is worth noting that our LIR produces better visual results that are more in line with the human aesthetic.
 The code is available at: https://github.com/Dongqi-Fan/LIR
-Keywords:
-Image Restoration Lightweight Attention
-[…]
-[width=1]figures/introduction.eps
-Figure 1: Visualization experiment of the image restoration process.
+Image Restoration; Patch Attention; Adaptive Filter; Residual Connection; Lightweight
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computational photographyccs: Computing methodologies Image processing
+Figure 1.
+Visualization experiment on the deraining task.
 (a) The paradigm of the most modern networks that incorporate both local and global residual connections.
-(b) The output by Restormer still exhibits visible traces of rain.
-(c) The output of our LIR.
-Which is more plain and solves the problem of residual connections carrying degradation.
-(d) The visualization of the three levels of the encoder and decoder based on Restormer.
-The top is deblurring and bottom is deraining.
-1 Introduction
+(b) The restored image output by the state-of-the-art model, Restormer (Zamir et al., 2022).
+It still exhibits visible traces of rain.
+(c) The restored image output by our LIR.
+It is cleaner and solves the problem of residual connections carrying degradation.
+(d) The visualization of intermediate feature maps of PReNet (Ren et al., 2019), SFNet (Cui et al., 2022), IRNeXt (Cui et al., 2023), Restormer, and LIR.
+1. Introduction
 Image Restoration (IR) is a fundamental computer vision task aiming to reconstruct a degraded image into a clean one, including deblurring, deraining, etc.
-In recent years, we have witnessed the great process of deep learning methods [28, 59, 52, 33, 39, 54, 67] in IR.
-However, many existing works only focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task.
+In recent years, we have witnessed the great process of deep learning methods (Liang et al., 2021; Zamir et al., 2021; Tu et al., 2022; Luo et al., 2020; Poirier-Ginter and Lalonde, 2023; Wang et al., 2022; Zhang et al., 2023c) in IR.
+However, many existing works tend to focus on designing the basic block and inefficiently stacking these blocks in a model, without considering the specific characteristics of the IR task.
 This approach often leads to parameter redundancy and computations unnecessary.
 Instead, our goal is an efficient and lightweight network design for various IR tasks.
-To address the issue of vanishing gradient and stabilize training, as shown in Fig.
-1 (a), most modern networks usually incorporate local and global residual connections.
-However, we think that these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output.
-To validate our idea, we conduct visualization experiments on the deblurring and deraining tasks using Restormer [58], a transformer model with a U-Net structure.
-Fig.
-1 illustrates the feature maps output by the encoder and decoder at three levels, as well as the input and output images of the Restormer, on the deblurring (at the top of Fig.
-1 (d)) and deraining (at the bottom of Fig.
-1 (d)) task.
-Notably, at the bottom of Fig.
-1 (d), it is evident that each feature map and the final output of the network (Fig.
-1 (b)) all contain degradation (rain).
-Although less pronounced in the deblurring task, a similar phenomenon can be observed at the top of Fig.
-1 (d).
-To overcome this issue, we use an ingenious structural design (transpose branchs in the LIR and LAA Fig.
-2) that effectively eliminates the degradation in the local and global residual connections.
-Fig.
-1 (c) shows the restored image output by our LIR, which is more plain than the Restormer (Fig.
-1 (b)).
-Moreover, we observe another serious issue in the visualization.
-In the visualization of the deblurring task (at the top of Fig.
-1 (d)), the feature maps generated by encoder_level2 and decoder_level2 exhibit clear contours while the other feature maps exhibit blurred contours.
-This issue is even worse in the deraining (at the bottom of Fig.
-1 (d)) task.
-All feature maps not only display very blurry contours but also contain a considerable amount of noise.
-Especially for the feature map generated by decoder_level1, it shows barely discernible contours, only with faint traces of rain.
+To address the issue of vanishing gradient and stabilize training, as shown in Figure 1 (a), most networks usually incorporate local and global residual connections.
+However, these connections may propagate the degradation of the input throughout the entire network and introduce it into the reconstructed clean output.
+To validate our idea, we conduct visualization experiments on deraining task using Restormer (Zamir et al., 2022), SFNet (Cui et al., 2022), IRNeXt (Cui et al., 2023), and PReNet (Ren et al., 2019).
+We select intermediate feature maps near the input and output of the networks for visualization because the semantics of the middle layer are too abstract.
+Results are shown in Figure 1 (d).
+Notably, it is evident that each feature map of these networks and the final output of the state-of-the-art model (Figure 1 (b)), Restormer, all contain severe degradation (rain).
+To overcome this issue, we use an ingenious structural design, using transpose branches in the LIR and LAA (Figure 2), that effectively eliminates the degradation in the local and global residual connections.
+The intermediate feature maps of our LIR (Figure 1 (d)) show better visual performance compared with the rest of the networks, and the restored image output by our LIR (Figure 1 (c)), which is more cleaner than the Restormer (Figure 1 (b)).
+More visual comparison of our LIR and these networks see Supplementary Materials.
+Moreover, we observe another issue in the visualization experiment.
+As shown Figure 1 (d)), the intermediate feature maps of the PReNet, SFNet, IRNetXt, and Resormer not only exhibit blurred contours but also contain a considerable amount of noise.
 The more important is that the high-frequency information has been lost.
-These observations indicate that although many modern networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and fails to address the core issues of image restoration, thus resulting in poor visual quality.
-Based on this insight, we propose the Adaptive Filter (Fig.
-3) (an component of the LAA) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
-Transformer-based models have demonstrated impressive performance in various tasks, both at high-level [12, 9, 31, 49] and low-level [6, 3, 54, 51, 27].
+These observations indicate that although many state-of-the-art networks have achieved satisfactory results on the metric (PSNR), the restoration process of the image is highly inefficient and missing to address the core issues of image restoration, thus resulting in poor visual quality.
+Based on this insight, we propose the Adaptive Filter (Figure 3) (a component of the LAA) as a solution to adaptively extract high-frequency information, enhance object contours, and remove degradation in various IR tasks.
+Transformer-based models have demonstrated impressive performance in various tasks, both at high-level (Dosovitskiy et al., 2020; Dong et al., 2022; Liu et al., 2021b; Touvron et al., 2021) and low-level (Chen et al., 2022b, 2021b; Wang et al., 2022; Tsai et al., 2022; Li et al., 2023).
 However, there exist different opinions.
-[10] argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs).
-In addition, [29, 47, 48] share a similar viewpoint that highlights the potential of MLPs beyond transformers.
-On the other hand, [50] suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy.
+(Dong et al., 2021) argues that attention is not all you need, emphasizing the importance of skip connections and Multi-layer Perceptrons (MLPs).
+In addition, (Liu et al., 2021a; Tolstikhin et al., 2021; Touvron et al., 2022) share a similar viewpoint that highlights the potential of MLPs beyond transformers.
+On the other hand, (Trockman and Kolter, 2022) suggests that CNN can achieve comparable results to transformers by leveraging the attention mechanism and patch embedding strategy.
 Thus, excessive computations can be avoided.
-Furthermore, [16, 32, 8, 70] take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer.
+Furthermore, (Guo et al., 2023; Liu et al., 2022; Ding et al., 2022; Zhou et al., 2022) take advantage of the attention mechanism in the CNN along with large convolution kernels to capture global information, instead of the self-attention among the transformer.
 Satisfactory results are achieved and excessive computations are avoided.
-Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module (Fig.
-4) (an component of the Attention Block) that aims to capture global information of the features and keeps parameters and computations friendly.
-In this paper, we propose a Lightweight Image Restoration network called LIR (Fig.
-2) to efficiently remove degradation and adapt to various IR tasks.
+Inspired by these perspectives, through leveraging the patch embedding strategy and MLP, we propose the Patch Attention module (Figure 4) (an component of the Attention Block) that aims to capture global information of the features and keeps parameters and computations friendly.
+In this paper, we propose a Lightweight Image Restoration network called LIR (Figure 2) to efficiently remove degradation and adapt to different IR tasks.
 The key component of LIR is the novel Lightweight Adaptive Attention Block (LAA) consisting of several proposed Adaptive Filters and Attention Blocks.
-Futhermore, the Attention Block is formed by proposed Patch Attention module (Fig.
-4), ResCABlock, and ResBlock (Details can be found in Fig.
-2).
-Thanks to the Adaptive Filter and Attention Block, the LAA Block is not only able to adaptively enhance the object contour and remove degradation but also to capture global information.
+Where the Attention Block is formed by proposed Patch Attention module (Figure 4), ResCABlock, and ResBlock (details can be found in Figure 2).
+Thanks to the Adaptive Filter and Attention Block, the LAA Block is capable of adaptively sharpening contours, removing degradation, and capturing global information in various Image Restoration scenes in a computation-friendly manner.
 Furthermore, we address the issue of degradation existing in the local and global residual connections by subtracting the output of the later component from the output of the first component in the LIR and LAA Block.
@@
 After that, we subtract this expanded degradation from the residual connection to obtain a clean residual connection.
-Extensive experiments demonstrate that LIR achieves comparable performance to state-of-the-art models in PSNR and produces better visual results that are more in line with the human aesthetic.
+Experiments demonstrate that our LIR achieves the state-of-the-art SSIM and comparable PSNR to state-of-the-art models on the deraining task, and comparable performance to state-of-the-art models with a parameter size of about 30% on the denoising, dehazing, and deblurring tasks.
 Our contributions can be summarized as follows:
-  • Through a simple structural design, the problem that local and global residual connections bring degradation is addressed, which is overlooked by modern networks.
-  • Given the problem of the loss of high-frequency information observed in the visualization experiments, we propose the Adaptive Filter (Fig.
-3) as a solution to adaptively extract high-frequency information and enhance object contours.
-  • To give LIR the ability to extract global information, we propose the Patch Attention module (Fig. 4), which is more computation-friendly than the self-attention of the transformer.
+  • Through an ingenious structural design, the problem that local and global residual connections bring degradation is addressed, which is overlooked by modern networks.
+  • Given the problem of the loss of high-frequency information observed in the visualization experiments, we propose the Adaptive Filter (Figure 3) as a solution to adaptively extract high-frequency information and enhance object contours.
+  • To give LIR the ability to extract global information, we propose the Patch Attention module (Figure 4), which is more computation-friendly than the self-attention of the transformer.
   • Based on Adaptive Filter, Patch Attention, ResCABlock, and ResBlock, we designed a lightweight and powerful block: LAA.
 It can make the LIR achieve comparable performance to state-of-the-art networks and better visual effects with smaller parameters and computations.
-2 Related Work
-2.1 Image Restoration
-In recent years, IR based on deep learning methods [13, 61, 55, 26, 63, 19] have shown promising results.
+2. Related Work
+Figure 2.
+The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Lightweight Adaptive Attention (LAA) Blocks.
+Thanks to the Adaptive Filter, Patch Attention, ResCABlock, and the ResBlock, the LAA Block is endowed with the ability to extract a variety of information efficiently.
+2.1. Image Restoration
+In recent years, IR based on deep learning methods (Du et al., 2020; Zhang et al., 2023a; Xie et al., 2021; Li et al., 2020; Zhang et al., 2022; Huang et al., 2023) have shown promising results.
 The transformer architectures achieve adaptive spatial aggregation at the cost of heavy computation.
-To address this challenge, a Kernel Basis Attention (KBA) module has been proposed by KBNet [67] in the CNN.
+To address this challenge, a Kernel Basis Attention (KBA) module has been proposed by KBNet (Zhang et al., 2023c) in the CNN.
 It adaptively aggregates spatial neighborhood information in a kernel-wise manner, thereby reducing computation.
-ADMS [38] proposes a Filter Attribution method based on Integral Gradient (FAIG) [55] to identify the contributions of the filters to remove specific degradations.
+ADMS (Park et al., 2023) proposes a Filter Attribution method based on Integral Gradient (FAIG) (Xie et al., 2021) to identify the contributions of the filters to remove specific degradations.
 By masking filters with minimal impact on degradation removal, the network’s efficiency in removing degradations is improved.
-IDR [61] proposes a learnable Principal Component Analysis (PCA) and views various IR tasks as multi-task learning methods to obtain network priors and remove degradations.
-MPRNet [59] introduces the Multi-Stage Progressive Restoration method, which is based on the U-Net architecture and follows a multi-input multi-output style.
-In addition, the contrastive learning paradigm is utilized by AirNet [24] to train the network to deal with different unknown degradations.
+IDR (Zhang et al., 2023a) proposes a learnable Principal Component Analysis (PCA) and views various IR tasks as multi-task learning methods to obtain network priors and remove degradations.
+MPRNet (Zamir et al., 2021) introduces the Multi-Stage Progressive Restoration method, which is based on the U-Net architecture and follows a multi-input multi-output style.
+In addition, the contrastive learning paradigm is utilized by AirNet (Li et al., 2022a) to train the network to deal with different unknown degradations.
 Through the prompting of the contrastive learning paradigm, the latent features in the network gradually away from the degradation space and close to the clean space.
-Different from the above works, our emphasis is on lightweight and efficient design while ensuring good performance.
-2.2 Vision Transformer for Image Restoration
-IPT [3] is the first work that introduces ViT [12] to low-level tasks to achieve image super-resolution, denoising, and deraining.
+The above works either involves a large number of parameters or a complex training process.
+Instead, our emphasis is on lightweight and efficient design while ensuring good performance.
+2.2. Vision Transformer for Image Restoration
+IPT (Chen et al., 2021b) is the first work that introduces ViT (Dosovitskiy et al., 2020) to low-level tasks to achieve image super-resolution, denoising, and deraining.
 The computational complexity of the self-attention part in ViT is $O(N^{2}D)$, where $N=H$$\times$$W$ represents the length of the patch sequence and $D$ represents the dimension of the sequence.
@@
 The computational complexity of the self-attention increases quadratically with the input size.
-To mitigate this issue, TransWeather [53] introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$.
-Uformer [54] and StripFormer [51] proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size.
-To further reduce the computational complexity, [22] estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL [27] proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention.
+To mitigate this issue, TransWeather (Valanarasu et al., 2022) introduces a reduction ratio factor R to control the complexity, resulting in a computational complexity of $O(N^{2}D/R)$.
+Uformer (Wang et al., 2022) and StripFormer (Tsai et al., 2022) proposes the window and stripe attention mechanisms respectively, which reduce the computational complexity to $O(NM^{2}D)$ and $O(D(H^{2}W+HW^{2}))$ by calculating dependencies in smaller windows and strips, where M represents the window size.
+To further reduce the computational complexity, (Kong et al., 2023) estimates the attention map that is output by the self-attention part using an element-wise product operation in the frequency domain, and GRL (Li et al., 2023) proposes an anchor stripe attention mechanism to futher narrow down the computational range of the slef-attention.
 The computational complexities are reduced to $O(NlogND)$ and $O(NMD)$ respectively, where M represents the size of feature map.
 However, the computation required for self-attention remains unaffordable.
-To address this issue, we propose a novel Patch Attention module to approximates the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation, leading to significant computational savings.
-[…]
-[width=]figures/model_2.eps
-Figure 2: The detailed structure of the proposed Lightweight Image Restoration (LIR) network, which is mainly stacked by Lightweight Adaptive Attention (LAA) Blocks.
-Thanks to the Adaptive Filter, Patch Attention, ResCABlock, and the ResBlock, the LAA Block is endowed with the ability to extract a variety of information efficiently.
-3 Method
-3.1 Lightweight Image Restoration
-The overall structure of LIR is illustrated in Fig.
-2, which primarily consists of Lightweight Adaptive Attention (LAA) Blocks.
+To address this issue, we propose a novel Patch Attention module to approximate the self-attention part by simplifying two matrix multiplications into one element-wise multiplication and removing the softmax operation for computational savings.
+3. Method
+Figure 3. The detailed architecture of the proposed Adaptive Filter. It is a reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
+Figure 4. The flow of the proposed Patch Attention.
+3.1. Lightweight Image Restoration
+The overall structure of LIR is illustrated in Figure 2, which primarily consists of Lightweight Adaptive Attention (LAA) Blocks.
 Both LIR and LAA Block adopt a down-up style to reduce computation and save parameters, that is downsizing the feature map in the beginning and upsizing at the end.
-To preserve low-frequency information and ensure training is stable, we add two global residual connections into LIR.
-One connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv.
-As mentioned in Sec.
-1, these residual connections can bring degradation.
+To preserve low-frequency information and ensure training is stable, we add two global residual connections and one local residual connection into LIR.
+For the global connections, one connects the input of the network to the output, while the other connects the head $3\times 3$ Conv to the tail $3\times 3$ Conv.
+For the local one, it connects the head and the tail.
+As mentioned in Section 1, these residual connections can bring degradation.
 To address this, we subtract the output of the fourth LAA Block from the output of the first LAA Block to obtain the corresponding degradation.
@@
 Next, we illustrate how we can achieve lightweight and efficiency through the Adaptive Filter and Attention Block in detail.
-[…]
-[width=]figures/filters_2.eps
-Figure 3: The comparison between our Adaptive Filter (left) and Edge-oriented Convolution Block (right).
-Both are reparameterization style that multiple convolutions are reparameterized into one during inference after the training.
-3.2 Adaptive Filter
-To achieve lightweight and efficiency, we first introduce the Adaptive Filter (at the left of Fig.
-3), inspired by Edge-oriented Convolution Block (ECB) [66] (at the right of Fig.
-3).
-The rich filters of different kinds allow the AF to enhance the object contour, remove degradation, and extract useful high-frequency information.
-Furthermore, the LAA block can be reparameterized as one $3\times 3$ and one high-pass filter after training to achieve lightweight.
-Although Adaptive Filter shares similarities with the ECB, it significantly outperforms ECB in three ways.
+3.2. Adaptive Filter
+To achieve lightweight and efficiency, we first introduce the Adaptive Filter (Figure 3), inspired by Edge-oriented Convolution Block (ECB) (Zhang et al., 2021b).
+The rich filters of different kinds allow the Adaptive Filter to enhance the object contour, remove degradation, and extract useful high-frequency information.
+Furthermore, the Adaptive Filter can be reparameterized as one $3\times 3$ and one high-pass filter after training to achieve lightweight.
+Figure 5. The visual comparisons for LIR, Restormer (Zamir et al., 2022), SFNet (Cui et al., 2022), and HINet (Chen et al., 2021a). LIR produces state-of-the-art visual results.
+Our Adaptive Filter significantly outperforms ECB in three ways.
 Firstly, the ECB introduces more convolution operations during training leading to more training costs (i.e.
@@
 Where Laplacian-4 and Laplacian-8 represent 4 domains and 8 domains; The high-pass filter obtains high-frequency information by subtracting the features after Gaussian blur processing from the original features.
-This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the AF for stages where the features have scarce noise and a clear outline.
-3.3 Attention Block
-While different types of filters in Adaptive Filter (AF) enable LAA blocks to extract rich information, AF alone is not enough, as we need to take full advantage of the rich information extracted by AF.
+This filter is more robust in noisy scenes compared to the above filters; Finally, we add a residual connection to the Adaptive Filter for stages where the features have scarce noise and a clear outline.
+3.3. Attention Block
+While different types of filters in Adaptive Filter enable LAA blocks to extract rich information, Adaptive Filter alone is not enough, as we need to take full advantage of the rich information extracted by Adaptive Filter.
 Thus, a strong and efficient Attention Block is introduced, which consists of four components: ResBlock, Patch Attention (PA), SimpleGate, and ResCABlock.
 We first use the ResBlock to preprocess input features, followed by the extraction of global information using our PA module.
-The efficacy of the gate function has been demonstrated in prior research [29, 52, 4].
-Here, we simply employ the SimpleGate module from [4] to further extract dependencies between features.
+The efficacy of the gate function has been demonstrated in prior research (Liu et al., 2021a; Tu et al., 2022; Chen et al., 2022a).
+Here, we simply employ the SimpleGate module from (Chen et al., 2022a) to further extract dependencies between features.
 Finally, we leverage the ResCABlock to extract channel-wise information.
-3.3.1 ResBlock and ResCABlock
+Table 1. Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset.
+Method Type Rain100L (Yang et al., 2017) Params [M] GFLOPs
+PSNR SSIM
+UMRL (Yasarla and Patel, 2019) CNN 29.18 0.923 0.98 9.31
+PReNet (Ren et al., 2019) RNN 32.44 0.950 0.17 37.27
+AirNet (Li et al., 2022a) CNN 34.90 0.966 8.93 169.47
+MPRNet (Zamir et al., 2021) CNN 36.40 0.965 20.13 960.39
+SPAIR (Purohit et al., 2021) CNN 36.93 0.969 - -
+HINet (Chen et al., 2021a) CNN 37.28 0.970 88.67 95.92
+IRNeXt (Cui et al., 2023) CNN 38.14 0.972 5.46 23.60
+MAXIM-2S (Tu et al., 2022) MLP 38.06 0.977 22.50 80.13
+SFNet (Cui et al., 2022) CNN 38.21 0.974 13.27 70.00
+DRCNet (Li et al., 2022b) CNN 38.23 0.976 18.90 -
+Restormer (Zamir et al., 2022) Transformer 38.99 0.978 26.13 92.85
+LIR (Ours) CNN 38.96 0.983 7.31 62.51
+3.3.1. ResBlock and ResCABlock
 Both ResBlock and ResCABlock are up-down style residual blocks that include a learnable parameter in the trunk.
 The only difference between ResBlock and ResCABlock is that the latter has a channel attention module at the tail.
-The Channel Attention module [68], based on Squeeze-and-Excitation (SE) [17], has gained much popularity for its ability to extract channel-wise information.
-Many studies [27, 6, 4, 69] have directly utilized this module in their models.
+The Channel Attention module (Zhang et al., 2018a), based on Squeeze-and-Excitation (SE) (Hu et al., 2018), has gained much popularity for its ability to extract channel-wise information.
+Many studies (Li et al., 2023; Chen et al., 2022b, a; Zhao et al., 2020) have directly utilized this module in their models.
 Following them, in our Attention Block, we also incorporate the Channel Attention module into the ResBlock to extract channel-wise information.
 In addition, the structure of ResBlock and ResCABlock are designed to be simple to meet lightweight and efficient requirements.
-[…]
-[width=]figures/patch_attention_2.eps
-Figure 4: The flow of the proposed Patch Attention.
-3.3.2 Patch Attention
-To give the Attention Block the ability to capture global information, we introduce the Patch Attention (PA) (Fig.
-4).
+Figure 6. The visual comparisons for LIR, Restormer (Zamir et al., 2022), AirNet (Li et al., 2022a), and MPRNet (Zamir et al., 2021). LIR produces state-of-the-art visual results.
+3.3.2. Patch Attention
+To give the Attention Block the ability to capture global information, we introduce the Patch Attention (PA) (Figure 4).
 In the self-attention part of the transformer, the attention map is calculated using two matrix multiplications and a softmax operation between q, k, and v, which incurs significant computational overhead.
@@
 The $(P_{h},P_{w})$ in the LIR is set to $(8,8)$, $(4,4)$, and $(2,2)$ respectively, and the $C_{scale}$ is set to 16 and 8, according to the depth of the network.
-Table 1: Image Deraining results evaluated by PSNR and SSIM on Rain100L dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
-[…]
-0.9 Method Type Rain100L [56] Params [M] GFLOPs PSNR $\downarrow$ SSIM UMRL [57] CNN 29.18 0.923 0.98 9.31 PReNet [42] RNN 32.44 0.950 0.17 37.27 AirNet [24] CNN 34.90 0.966 8.93 169.47 MPRNet [59] CNN 36.40 0.965 20.13 960.39 SPAIR [40] CNN 36.93 0.969 $\triangle$ $\triangle$ HINet [5] CNN 37.28 0.970 88.67 95.92 MAXIM-2S [52] MLP 38.06 0.977 22.50 80.13 SFNet [7] CNN 38.21 0.974 13.27 70.00 DRCNet [25] CNN 38.23 0.976 18.90 $\triangle$ Restormer [58] Transformer 38.99 0.978 26.13 79.31 LIR (Ours) CNN 38.96 0.983 7.31 63.96
-[…]
-[width=]figures/derain100l.eps
-Figure 5: The visual comparisons for LIR, Restormer [58], SFNet [7], and HINet [5]. LIR produces state-of-the-art visual results.
-4 Experiment
-4.1 Experiment Settings
+Figure 7. The visual comparisons for LIR, FFA-Net (Qin et al., 2020), GridDehazeNet (Liu et al., 2019), and AirNet (Li et al., 2022a). LIR produces better visual results.
+Table 2. Image denoising results evaluated by PSNR on CBSD68 and Urban100 datasets. Where the background of light blue for the method that is learned to handle various noise levels.
+Method Type CBSD68 (Martin et al., 2001) Urban100 (Huang et al., 2015) Params [M] GFLOPs
+$\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$
+FD-GAN (Dong et al., 2020a) GAN 30.25 28.81 26.43 - - - 13.98 38.90
+DL (Fan et al., 2019) CNN 33.05 30.41 26.90 - - - - -
+MPRNet (Zamir et al., 2021) CNN 33.54 30.89 27.56 - - - 20.13 960.39
+FFDNet (Zhang et al., 2018b) CNN 33.87 31.21 27.96 33.83 31.40 28.05 - -
+DnCNN (Zhang et al., 2017) CNN 33.90 31.24 27.95 32.98 30.81 27.59 - -
+AirNet (Li et al., 2022a) CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47
+BRDNet (Tian et al., 2020) CNN 34.10 31.43 28.16 34.42 31.99 28.56 - -
+IDR (Zhang et al., 2023b) CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 -
+DRUNet (Zhang et al., 2021a) CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78
+Restormer (Chen et al., 2022a) Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 92.85
+LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 62.51
+4. Experiment
+4.1. Experiment Settings
 We set the number of LAA Blocks in the LIR to [3,3,3,3,4], and the width to 48.
@@
 Random image cropping and horizontal flipping are used as data augmentation strategies.
-Following [58], we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising.
+Following (Zamir et al., 2022), we use the cosine annealing strategy to gradually decrease the learning rate from $3e^{-4}$ to $1e^{-6}$, and use the progressive learning, where training patch size and batch size pairs are: [($128^{2}$,32), ($160^{2}$,16), ($256^{2}$,8)] at iterations [250K, 250K, 250K] for deblurring, deraining, and denoising.
 For dehazing, we set [($128^{2}$,32), ($160^{2}$,16)] at iterations [350K, 400K].
@@
 Specifically, parameters are counted using the PyTorch, and the computation is measured using the thop package.
-All experiments are conducted using the official release code for each method.
-For methods that do not have released code or are not based on PyTorch, we have marked them with $\triangle$ in the tables.
-[…]
-[width=0.6]figures/dehaze.eps
-Figure 6: The visual comparisons for LIR, FFA-Net [41], GridDehazeNet [30], and AirNet [24]. LIR produces better visual results.
-4.2 Image Deraining
-The Rain14000 [15] (11,200 training images) and Rain100H [56] (1,254 training images) are used to train LIR, and the Rain100L [56] datasets is used for evaluation.
-The results are presented in Tab.
-1.
-In the Rain100L results (Tab.
-1), LIR with minimal parameters and computation outperforms the DRCNet [25] and achieves comparable performance to Restormer [58] with only a slight 0.02 dB decrease on PSNR.
+4.2. Image Deraining
+The Rain14000 (Fu et al., 2017) (11,200 training images) and Rain100H (Yang et al., 2017) (1,254 training images) are used to train LIR, and the Rain100L (Yang et al., 2017) datasets is used for evaluation.
+The results are presented in Table 1, where LIR outperforms the DRCNet (Li et al., 2022b) and achieves comparable performance to Restormer (Zamir et al., 2022) with only a slight 0.02 dB decrease on PSNR with minimal parameters.
 In addition, our LIR achieves state-of-the-art on SSIM, which significantly outperforms Restormer.
-As GFLOPs do not accurately reflect the computations of the transformer model, the actual computation of LIR is much lower compared to Restormer, as shown in Tab.
-1.
-It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet [7], the same lightweight design method as ours.
-Fig.
-5 shows visual comparisons among LIR, Restormer, SFNet, and HINet.
+It is worth noting that with fewer parameters and computations, LIR significantly outperforms SFNet (Cui et al., 2022), the same lightweight design method as ours.
+Figure 5 shows visual comparisons among LIR, Restormer, SFNet, and HINet.
 It can be demonstrated that LIR produces state-of-the-art visual results, despite LIR having a slightly lower PSNR than Restormer.
-Table 2: Image dehazing results evaluated by PSNR and SSIM on SOTS outdoor dataset. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
-[…]
-0.9 Method Type SOTS outdoor [23] Params [M] GFLOPs PSNR $\downarrow$ SSIM FD-GAN [11] GAN 23.15 0.920 13.98 38.90 AirNet [24] CNN 23.18 0.900 8.93 169.47 IDR [62] CNN 25.24 0.943 15.34 $\triangle$ GridDehazeNet [30] CNN 30.86 0.982 0.96 12.05 U2-Former [20] Transformer 31.10 0.976 $\triangle$ $\triangle$ FFA-Net [41] CNN 33.38 0.980 4.46 161.74 MAXIM-2S [52] MLP 34.19 0.985 22.5 80.13 DehazeFormer [44] Transformer 34.95 0.984 2.52 13.37 LIR (Ours) CNN 34.85 0.985 7.31 63.96
-[…]
-[width=]figures/denoise.eps
-Figure 7: The visual comparisons for LIR, Restormer [58], AirNet [24], and MPRNet [59]. LIR produces state-of-the-art visual results.
-4.3 Image Dehazing
-The OTS [23] (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor [23] is used for evaluation.
-In the SOTS outdoor results (Tab.
-2), LIR significantly outperforms both the MAXIM-2S [52] (MLP-based) and the U2-Former [20] (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB.
+Table 3. Image dehazing results evaluated by PSNR and SSIM on SOTS outdoor dataset.
+Method Type SOTS outdoor (Li et al., 2018) Params [M] GFLOPs
+PSNR SSIM
+FD-GAN (Dong et al., 2020a) GAN 23.15 0.920 13.98 38.90
+AirNet (Li et al., 2022a) CNN 23.18 0.900 8.93 169.47
+IDR (Zhang et al., 2023b) CNN 25.24 0.943 15.34 -
+GCANet (Chen et al., 2019) CNN 30.23 0.980 - -
+GridDehazeNet (Liu et al., 2019) CNN 30.86 0.982 0.96 13.69
+U2-Former (Ji et al., 2021) Transformer 31.10 0.976 - -
+FFA-Net (Qin et al., 2020) CNN 33.38 0.980 4.46 161.74
+MSBDN (Dong et al., 2020b) CNN 33.67 0.985 - -
+MAXIM-2S (Tu et al., 2022) MLP 34.19 0.985 22.5 80.13
+LIR (Ours) CNN 34.85 0.985 7.31 62.51
+Table 4. Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets.
+Method Type GoPro (Nah et al., 2017) HIDE (Shen et al., 2019) Params [M] GFLOPs
+PSNR SSIM PSNR SSIM
+IDR (Zhang et al., 2023b) CNN 27.87 0.846 - - 15.34 -
+DeblurGAN-v2 (Kupyn et al., 2019) GAN 29.55 0.934 26.61 0.875 59.16 20.12
+SRN (Tao et al., 2018) RNN 30.26 0.934 28.36 0.915 - -
+Gao et al. (Gao et al., 2019) CNN 30.90 0.935 29.11 0.913 - -
+DBGAN (Zhang et al., 2020) GAN 31.10 0.942 28.94 0.915 11.58 189.71
+MT-RNN (Park et al., 2020) RNN 31.15 0.945 29.15 0.918 - -
+DMPHN (Zhang et al., 2019) CNN 31.20 0.940 29.09 0.924 21.70 -
+Zhe et al. (Jiang et al., 2020) CNN 31.79 0.949 - - - -
+Suin et al. (Suin et al., 2020) CNN 31.85 0.948 29.98 0.939 - -
+SPAIR (Purohit et al., 2021) CNN 32.06 0.950 30.29 0.931 - -
+LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 62.51
+4.3. Image Dehazing
+The OTS (Li et al., 2018) (72,315 training images) is used to train LIR, and the Synthetic Objective Testing Set (SOTS) outdoor (Li et al., 2018) is used for evaluation.
+In the SOTS outdoor results (Table 3), LIR significantly outperforms both the MAXIM-2S (Tu et al., 2022) (MLP-based) and the U2-Former (Ji et al., 2021) (transformer-based) with improvement in PSNR of 0.66 dB and 3.75 dB.
 In addition, LIR achieves comparable performance on PSNR and better performance on SSIM to DehazeFormer.
-It is worth noting that GFLOPs and parameters do not accurately reflect the complexity of the transformer model, the actual complexity of LIR is much lower compared to DehazeFormer.
-Visual comparisons are shown in the Fig.
-6.
-Table 3: Image denoising results evaluated by PSNR and SSIM on CBSD68 and Urban100 datasets.
-Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch and light blue for the method that is learned to handle various noise levels.
-[…]
-0.9 Method Type CBSD68 [35] $\downarrow$ Urban100 [18] Params [M] GFLOPs $\sigma=15$ $\sigma=25$ $\sigma=50$ $\sigma=15$ $\sigma=25$ $\sigma=50$ FD-GAN [11] GAN 30.25 28.81 26.43 - - - 13.98 38.90 DL [14] CNN 33.05 30.41 26.90 - - - $\triangle$ $\triangle$ MPRNet [59] CNN 33.54 30.89 27.56 - - - 20.13 960.39 AirNet [24] CNN 33.92 31.26 28.00 34.40 32.10 28.88 8.93 169.47 BRDNet [46] CNN 34.10 31.43 28.16 34.42 31.99 28.56 $\triangle$ $\triangle$ IDR [62] CNN 34.11 31.60 28.14 33.82 31.29 28.07 15.34 $\triangle$ DRUNet [64] CNN 34.30 31.69 28.51 34.81 32.60 29.61 32.64 80.78 Restormer [4] Transformer 34.39 31.78 28.59 35.06 32.91 30.02 26.13 79.31 LIR (Ours) CNN 34.31 31.69 28.50 34.81 32.59 29.57 7.31 63.96
-4.4 Image Denoising
+Visual comparisons are shown in the Figure 7.
+4.4. Image Denoising
 We perform denoising experiments on synthetic benchmark datasets generated by additive white Gaussian noise.
-The DFWB (DIV2K [1], Flickr2K, WED [34], BSD500 [2]) datasets with 8,594 training images are used to train LIR.
-For evaluation, we use the CBSD68 [35] and Urban100 [18] datasets.
-The light blue color in the Tab.
-3 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50).
+The DFWB (DIV2K (Agustsson and Timofte, 2017), Flickr2K, WED (Ma et al., 2016), BSD500 (Arbelaez et al., 2010)) datasets with 8,594 training images are used to train LIR.
+For evaluation, we use the CBSD68 (Martin et al., 2001) and Urban100 (Huang et al., 2015) datasets.
+The background of the light blue color in the Table 2 indicates that the method is trained using a blind denoising setting (random $\sigma$ from [0,50]) and then evaluated under a fixed $\sigma$ (0, 25, or 50).
 Other methods are trained with a fixed $\sigma$ and evaluated with the same $\sigma$ as training.
-As shown in the Tab.
-3, LIR achieves comparable performance to Restormer [58] and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting.
-Moreover, it is worth noting that with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet [64] on Urban100, while the DRUNet is trained under fixed $\sigma$ setting and our LIR is under random $\sigma$ setting.
-The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Fig.
-7.
-Although LIR is not as high as Restormer on PSNR, the state-of-the-art visual result is achieved.
-Table 4: Image deblurring results evaluated by PSNR and SSIM on GoPro and HIDE datasets. Where $\triangle$ stands for the official code that is unavailable or not based on PyTorch.
-[…]
-0.9 Method Type GoPro [36] HIDE [43] Params [M] GFLOPs PSNR $\downarrow$ SSIM PSNR $\downarrow$ SSIM IDR [62] CNN 27.87 0.846 - - 15.34 $\triangle$ DBGAN [65] GAN 31.10 0.942 28.94 0.915 11.58 189.71 MT-RNN [37] RNN 31.15 0.945 29.15 0.918 $\triangle$ $\triangle$ DMPHN [60] CNN 31.20 0.940 29.09 0.924 21.70 $\triangle$ Zhe et al.
-[21] CNN 31.79 0.949 - - $\triangle$ $\triangle$ Suin et al.
-[45] CNN 31.85 0.948 29.98 0.939 $\triangle$ $\triangle$ SPAIR [40] CNN 32.06 0.950 30.29 0.931 $\triangle$ $\triangle$ LIR (Ours) CNN 32.19 0.931 30.31 0.917 7.31 63.96
-4.5 Image Deblurring
-The GoPro [36] (2,103 training images) dataset is used to train LIR, and the GoPro (1,111 images) and HIDE [43] (2,025 images) datasets are used for evaluation.
-Tab.
-4 demonstrates that our LIR significantly outperforms other methods with fewer parameters and computations.
-Specifically, LIR is 0.13 dB higher than SPAIR [40], and 0.34 dB significantly higher than Suin et al [45].
-4.6 Ablation Studies
-We conduct ablation experiments on the Rain100L dataset, using unchanged LIR as the baseline.
-The ablation experiments are performed on three types: Simple Residual, No Patch Attention, and No Adaptive Filter.
-Simple Residual refers to using vanilla global and local residual connections in the model, without removing the degradation in residual connections through structural design.
-In addition, No Adaptive Filter refers to using a vanilla convolution to replace the Adaptive Filter for fair comparison.
-Therefore, the number of parameters is unchanged.
-The results in the Tab.
-5 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR.
-In addition, the experiment without Patch Attention demonstrates that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
-Table 5: Ablation stuides are performed on Rain100L dataset.
-Metric Baseline Simple Residual Patch Attention Adaptive Filter
-PSNR 38.96 38.87 38.85 38.66
-Params [M] 7.31 6.86 6.51 7.31
-GFLOPs 63.96 49.15 62.15 62.40
-5 Conclusion and Limitation
-5.1 Conclusion
+As shown in the Table 2, LIR achieves comparable performance to Restormer (Zamir et al., 2022) and significantly outperforms all other models on the CBSD68 dataset with a blind denoising setting.
+Moreover, it is worth noting that with a 77% reduction in parameters and a 20% reduction in computations, our LIR achieves comparable performance to DRUNet (Zhang et al., 2021a) on Urban100, while the DRUNet is trained under fixed $\sigma$ setting and our LIR is under random $\sigma$ setting.
+The visual comparisons between LIR, Restormer, AirNet, and MPRNet are shown in Figure 6.
+Although LIR is not as high as Restormer on PSNR, the better visual result is achieved.
+4.5. Image Deblurring
+The GoPro (Nah et al., 2017) (2,103 training images) dataset is used to train LIR, and the GoPro (1,111 images) and HIDE (Shen et al., 2019) (2,025 images) datasets are used for evaluation.
+Table 4 demonstrates that our LIR significantly outperforms other methods with fewer parameters and computations.
+Specifically, LIR is 0.13 dB higher than SPAIR (Purohit et al., 2021), and 0.34 dB significantly higher than Suin et al (Suin et al., 2020).
+4.6. Ablation Studies
+We conduct ablation experiments on the deraining task using different types.
+Type 1 refers to LIR, Type 2 refers to using vanilla global and local residual connections in the model without removing the degradation in residual connections through structural design, Type 3 refers to no Patch Attention module, Type 4 refers to no Adaptive Filter, and Type 5 refers to using vanilla global and local residual connections in the model and not include Patch Attention module and Adaptive Filter.
+The results in the Table 5 demonstrate that the absence of the Adaptive Filter has the greatest negative impact on LIR, where PSNR is droped by 0.3 dB.
+Furthermore, it also shown that our proposed Patch Attention module is efficient and computation-friendly, only bringing 0.8 M parameters and 1.81 GFLOPs computations.
+Although Type 1 only improved performance by 0.09 dB compared to Type 2, structural design for residual connections is the key to the final visual improvement.
+In addition, we present the intermediate feature maps of Type 1 and Type 4 (Figure 9).
+Obviously, LIR with Adaptive Filter has a better visual result than without Adaptive Filter, with less noise and rain trails.
+We also present the intermediate feature maps of Type 1 and Type 2 (Figure 8), where the former has a better visual result than the latter, with fewer rain trails and better contour.
+More visual results about Type 1, Type 2, and Type 5 please see Supplementary Materials.
+Table 5.
+Ablation studies are performed on Rain100L dataset.
+Where Vanilla Con.
+(Connections) refers to using vanilla global and local residual connections in the model without removing the degradation in residual connections.
+Type 1 Type 2 Type 3 Type 4 Type 5
+Vanilla Con.
+Adaptive Filter
+Patch Attention
+Params [M] 7.31 6.86 7.31 6.51 5.88
+GFLOPs 62.51 49.15 62.51 62.15 49.01
+PSNR 38.96 38.87 38.66 38.85 37.65
+5. Conclusion and Limitation
+Figure 8. The visualization of intermediate feature maps of the Type 1 (top) and the Type 2 (bottom) in the deraining task.
+Figure 9. The visualization of intermediate feature maps of the Type 1 (top) and the Type 4 (bottom) in the deraining task.
+5.1. Conclusion
 In this paper, we introduce a Lightweight Image Restoration network called LIR, which aims to efficiently remove degradation.
-Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and a simple structural design.
-In addition, we propose the Adaptive Filter (AF) to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks.
-Our AF achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Tab.
-5).
-To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention (PA) module by leveraging an MLP and patch embedding strategy.
-5.2 Limitation
-For the deraining task, LIR achieves comparable performance with the state-of-the-art method on PSNR and the best performance on SSIM with smaller parameter numbers and computations; For denoising and dehazing tasks, LIR also achieves comparable performance to the state-of-art method.
-In addition, LIR achieves state-of-the-art visual results on deraining and denoising tasks.
-However, as Tab.
-4 shows, LIR is less prominent on deblurring tasks than deraining and denoising.
+Firstly, the problem that local and global residual connections propagate the degradation throughout the entire network is addressed through the use of the Transpose operation and an ingenious structural design.
+In addition, we propose the Adaptive Filter to adaptively extract high-frequency information, enhance object contours, and eliminate degradation in various IR tasks.
+Our Adaptive Filter achieves promising results while requiring minimal parameters and computations, as demonstrated in our ablation studies (Section 5).
+To capture global information while keeping small parameters and low computational requirements, we introduce the Patch Attention module by leveraging an MLP and patch embedding strategy.
+On the deraining task, our LIR achieves the state-of-the-art Structure Similarity Index Measure and comparable performance to state-of-the-art models on Peak Signal-to-Noise Ratio.
+For denoising, dehazing, and deblurring tasks, LIR also achieves a comparable performance to state-of-the-art models with only about 30% size.
+5.2. Limitation
+LIR achieves good performance on deraining, denoising and dehazing tasks and better visual results.
+However, as Figure 4 shows, LIR is less prominent on deblurring tasks than deraining and denoising.
 We believe that this is because, in the deblurring task, object outlines tend to be stretched out very long.